P = 50-0.1Q, P=20$

Please kindly explain about this exercise.


1
Expert's answer
2021-01-01T14:01:38-0500

"20 = 50-0.1Q \\

20-50 = -0.1Q \\

-30 = -0.1Q \\

30 = 0.1Q \\

10 \times 30 = 10 \times 0.1Q \\

300 = Q \\

Q = 300"